OpenStudy (anonymous):

A rectangular park 60mlong and 40m wide has two concrete crossroads running in the middle of the park and the rest of the park is used as a lawn .if the area of the lawn is 2109 sq.m.then what is the width of the road?

OpenStudy (mandre):

 Area of park = 2400 Area of roads = 2400 - 2109 = 291 40x + 60x - x^2 = 291 -x^2 + 100x - 291 = 0